Fermi National Laboratory
 

Accelerator Update for April 16, 2007 - April 18, 2007

 

Monday April 16

The day shift began with Operations monitoring Tevatron (TeV) store 5355 (187.88E30) and the Recycler stash (111.5E10), with the Antiproton Source (Pbar) stacking (13.5E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams. Interlock technicians were conducting safety system testing in Switchyard (SY).

The Neutron Therapy Facility (NTF) reported at 10:38 AM that they had completed treating patients for the day.

MI experts began a stacktail study at 11:13 AM.

The I- Source vacuum gauge tripped off at 11:17 AM. It reset.

Operations turned off Linac beam at 2:11 PM so that experts could begin repairs on a Klystron RF station (KRF4).

Linac experts completed the KRF4 work. Linac resumed sending beam at 2:58 PM.

At 4:18 PM, LRF3 tripped off and wouldn’t reset. Linac experts investigated and reset the station.

Operations reported at 4:25 PM that the Recycler had stashed antiprotons twice during the day shift.

The MI studies ended at 5:07 PM.

At 10:51 PM, Operations reported that the stash aborted and many other devices alarmed. The store and the stack remained intact. Operators investigated and eventually determined that there was a power glitch.

Operations prepared MI for an access at 11:41 PM so that experts could repair the Recycler dampers.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Tuesday April 17

The midnight shift began with Operations monitoring store 5355 (55.38E30) and the stack (63.0E10), with no stash to monitor, and with MI in access to repair the Recycler dampers. Pbar, NuMI, and MiniBooNE were off due to the MI access.

At 12:29 AM, BRF14 tripped off for the first of five times in three minutes. Operators left the station off after the last trip so it could rest.

The Recycler damper work ended at 1:18 AM.

Operations tried to turn BRF14 back on at 1:26 AM, but it wouldn’t stay on. Operators left the station off so it could rest a little longer.

Operations established beam to MiniBooNE at 2:09 AM.

Operations established beam to Pbar and NuMI at 2:30 PM.

At 3:22 AM, Operations reported that a fire alarm had caused the Crew Chief to turn off all beam and that everyone abandoned the MCR for about 10 minutes. The Fire Chief reported that there was a sprinkler pressure surge. The Linac resumed sending beam at 3:33 AM.

At 6:07 AM, Operators again tried and failed to turn on BRF14. This time they bypassed the station and turned on BRF19 in its place.

Operations reported at 7:58 AM that the Recycler had stashed twice during the midnight shift.

An RF expert turned on BRF14 at 8:11 AM. Operators turned off BRF19.

At 8:22 AM, Operations reported that the Accumulator BPMs at AP30 were not communicating properly. Operators contacted experts.

A controls expert restarted the AP30 BPMs at 8:57 AM.

At 10:08 AM, the Pelletron tripped off. Experts reported that they were loading data that shouldn’t have caused any problems, but during the load a device hung up (IRM). The experts will access their pit to reset the device.

Operations terminated store 5355 at 11:04 AM.

Operations prepared CDF and D0 for access at 11:23 AM. Both experiments were out of access by 12:22 PM.

Operations put the TeV through a dry squeeze at 12:57 PM.

Operations put the TeV into shot setup at 1:30 PM.

Operations established store 5357 at 3:44 PM with an initial luminosity of 163.33E30.

A TeV expert began a Tevatron Electron Lens (TEL1) study at 4 PM.

At 4:45 PM, LRF3 tripped off. It reset.

At 4:49 PM, Operations reported that the MI beam permit was tripping off due to increased beam intensity on the MI-8 collimators.

Pelletron studies ended at 5:07 PM.

At 6:22 PM, LRF3 tripped off. It reset.

A MI expert began a short barrier bucket study at 6:33 PM.

At 9:51 PM, Operations lowered the Booster intensity to control the MI beam permit trips.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Wednesday April 18

The midnight shift began with Operations monitoring store 5357 (65.57E30) and the stash (164.9E10), with Pbar stacking (12.6E10), and with NuMI and MiniBooNE taking beam.

At 3:33 AM, Operations reported a number of MI beam permit trips due to MI-8. They all reset.

At 3:41 AM, LRF3 tripped off. It reset.

At 3:49 AM, MI-8 caused MI aborts and caused the Pelletron to trip off. Everything reset.

At 4:10 AM, a fire alarm required the Crew Chief to turn off the Linac beam and for the crew to leave the MCR. The operators were back in the MCR and delivering beam by 4:27 AM. The Fire Chief reported another sprinkler pressure surge.

At 4:44 AM, LRF3 tripped off. It reset.

At 7:32 AM, NuMI tripped off due to the current on a quadrupole power supply drifting a bit low. Everything reset.

Operations reported at 7:50 AM that the Recycler had stashed three times during the midnight shift.

Operations terminated store 5357 at 8 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow experts to conduct various machine studies that don’t affect stacking or interfere with each other.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/18/2007    email Fermilab